Respected sir,
plz help me to solve this maths problem.
1] A certain farmaer pays $ 30 per acer per month to rent farmland.how much does the farmer pay per month to rent a rectangular plot of farmland tha is 360 feet by 605 feet? ( 43,560 sq.feet=1 acre)
waiting for your reply sir.
thank u.
regards
Binita.

Answer by nerdybill(7384)

A certain farmaer pays $ 30 per acer per month to rent farmland.how much does the farmer pay per month to rent a rectangular plot of farmland tha is 360 feet by 605 feet? ( 43,560 sq.feet=1 acre)
.
The area of the farmland is:
360*605 = 217800 square feet
.
Convert the above into acres:
217800 * 1/43560 = 5 acres
.
Rent, therefore, is:
5 * 30 = $150 per month